Which of the following is the best estimate of the circumference of a circle that has a diameter of 7 feet?
11 ft.
14 ft.
42 ft.
21 ft.

Answer by richwmiller(17219)

circumference=2*pi*r
diameter =2*r
circumference =d*pi
pi is about 3.14
3*7=21